10.00 points Consider the following information for Maynor Company, which uses a periodic inventory system Transaction Units Unit Cost Total Cost January 1 Beginning Inventory 31 March 28 Purchase August 22 Purchase October 14 Purchase 41 62 67 $81 87 91 97 2,511 3,567 5,642 6,499 Goods Available for Sale 201 $18,219 The company sold 67 units on May 1 and 62 units on October 28 Required Calculate the companys ending inventory and cost of goods sold using the each of follawing inventory costing methods. (Round the per unit cost to two decimal places and then round your answer to the nearest whole dollar.) a. Weighted Average Ending Inventory Cost of Goods Sold

Taking exam, please help! calculate ending inventory and cogs using the following inventory costing methods. THANKYOU
Answer #1

Calculate following :

Average cost per unit = 18219/201 = 90.64 per unit

Ending inventory = 72*90.64 = 6526

Cost of goods sold = 129*90.64 = 11693